OpenStudy (anonymous):

what is the range of h(x)=x+14

OpenStudy (anonymous):

all real numbers

OpenStudy (anonymous):

it is a line, goes from \(-\infty\) to \(\infty\)

OpenStudy (anonymous):

it s not just x its \[\sqrt{x}\]

OpenStudy (anonymous):

then the answer is different

OpenStudy (anonymous):

\[f(x)=\sqrt x+14\] ??

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

in that case the range is \([14,\infty)\) since \(\sqrt{x}\geq 0\) for any \(x\)

OpenStudy (anonymous):

do you know what would be the domain